OpenStudy (anonymous):

How many solutions exist for the system y = 2x + 5 and y = -0.2x - 4? no solution one solution infinitely many solutions its only one right

OpenStudy (frostbite):

2 equations 2 veriables.... trid to solve it?

OpenStudy (frostbite):

or you don't even have to....

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

yes so they can only intercect once right?

OpenStudy (frostbite):

Exactly.

OpenStudy (anonymous):

oh thanks

OpenStudy (anonymous):

how azbout this? How many solutions exist for the system, x + y = 1 and y = -x + 1? no solution one solution infinitely many solutions

OpenStudy (frostbite):

unless y=m_1x+b and y=m_2x+b where m_1=m_2 then they are what? ;)

OpenStudy (anonymous):

im sorry you lost me

OpenStudy (frostbite):

\[y=1-x=-x+1\]\[x+y=1\]

OpenStudy (frostbite):

sorry wrong order... but you see the rewriting?